If 11 students offer Biology, 13 students offer Maths, 14 students offer Agric, 9 students offer Maths and Biology, 3 students offer Maths and Agric, 7 students offer Maths, Biology and Agric. How many students offer Biology and Agric?


1
Expert's answer
2021-02-24T06:57:25-0500

Let the total number of students be "S" and the numbers of student offering Biology and Maths be "x" .

"S=n(B)+n(M)+n(A)+n(B\bigcap M)+n(B \bigcap A)+n(M\bigcap A)+n(B\bigcap M\bigcap A)\\
S=11+13+14+x+9+3+7\\
S=57+x\\
x=S-57"